OpenStudy (anonymous):

[4.02] What is the x-intercept and y-intercept of the graph of 3x - 4y = 12? x-intercept = (-4, 0); y-intercept = (0, 3) x-intercept = (4, 0); y-intercept = (0, -3) x-intercept = (0, -3); y-intercept = (4, 0) x-intercept = (0, 3); y-intercept = (-4, 0)

OpenStudy (anonymous):

when you let x=0 what do you get. When you let y=0 what do you gte?

OpenStudy (anonymous):

for x=0 3(0)-4y=12 -4y=12 for y = 0 3x-4(0)=12 3x=12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so it would be "B" x-intercept = (4, 0); y-intercept = (0, -3)